OpenStudy (anonymous):

select the best possible first step to solving the following system by first eliminating the y variable. 4x+5y=-8 3x+2y=1

hero (hero):

I'd solve for y first.

hero (hero):

for both equations

hero (hero):

That, to me, would be the way

hero (hero):

so you would have... y = -(8+4x)/5 and y = (1-3x)/2 Set y = y and then you're left with: -(8+4x)/5 = (1-3x)/2

hero (hero):

Then cross multiply and finish solving for x